OpenStudy (anonymous):

Solve: 2x + y = 7 x – y = –4

OpenStudy (anonymous):

x=4+y then insert into 2x+y=7

OpenStudy (anonymous):

oh sorry x=1 1-y=-4 y=5

OpenStudy (anonymous):